A car will go 3/8 of a given distance in one hour. What part will it cover in 5/8 of an hour?
Found 2 solutions by Jeff Gordon, dabanfield:
Answer by Jeff Gordon(65)

three eights of the given distance like you said earlier.

Answer by dabanfield(803)

A car will go 3/8 of a given distance in one hour. What part will it cover in 5/8 of an hour?
Let D be the GIVEN distance and R the rate. Then since rate = distance/time we have:
R = (3/8*D)/1
R = 3/8*D
Since distance = rate*time the distance d traveled in 5/8 hour is:
d = (3/8*D)*5/8
d = 15/64*D
So the part of the distance traveled is 15/64.